OpenStudy (anonymous):

evaluate the integral ∫ (4dx)/(3+6x)

OpenStudy (callisto):

\[\int \frac{4dx}{3+6x}\]Factorize the denominator, we get \[=\int\frac{4dx}{3(1+2x)}\]Pull out the constant 4/3, we get\[=\frac{4}{3}\int\frac{1}{1+2x}dx\]Do you know how to integrate it now?

OpenStudy (anonymous):

no sorry i dont know how

OpenStudy (callisto):

Do you know how to integrate \(\frac{1}{x}\) ?

OpenStudy (anonymous):

i think i've got it now 2/3 ln abs (-3-6x) +c

OpenStudy (callisto):

The abs(-3-6x) part does not seem right.

OpenStudy (anonymous):

abs(3+6x)